OpenStudy (anonymous):

Identify the third term in the expansion of (7x – 2y)^3

OpenStudy (unklerhaukus):

sub in a=7x, b=2y.

OpenStudy (unklerhaukus):

(a−b)^3=(a−b)(a−b)(a−b) =(a−b)(a^2−2ba+b^3) =(a^3−a^2b−2a^2b+2ab^2+ab^3−b^3)